OpenStudy (anonymous):

The sale price of a volley at a discount of 40% is $18. What was its original price?

OpenStudy (anonymous):

$45?

OpenStudy (anonymous):

nope 30

OpenStudy (anonymous):

30 is the answer

OpenStudy (anonymous):

you could check for yourself

OpenStudy (anonymous):

40 percent of 30 is 12 and 30-12=18 so 30 is the answer :D

OpenStudy (anonymous):

.6x=18 x=18/.6

OpenStudy (anonymous):

18/.6 is 30

OpenStudy (anonymous):

QWhere you get .6?

OpenStudy (anonymous):

(1-.4)=.6

OpenStudy (anonymous):

so let x be the original price of the product, now we know that if we discount. that means we subtract 40 percent of the original price, we will have 18. so if we write these in an equation. x-40%x=18 now 40 per cent means 40/100 (cent is 100 right?) 40%=40/100=.4 so: x-.4x=18 by distributive property x(1-.4)=18 now 1-.4 is .6 so x(.6)=18 x=18/.6 x=30 :D